A Carnot engine on a ship extracts heat from seawater at 16.0∘C and exhausts the heat to evaporating dry ice at -78.0∘C.

1) If the ship's engines are to run at 7400 horsepower, what is the minimum amount of dry ice the ship must carry for the ship to run for a single day? Use that 1 horsepower = 746 W. The latent heat of sublimation of carbon dioxide is Ls = 573,700 J/kg. (Express your answer to three significant figures.)
Answer #1

Energy required for a day E= 24*3600*7400*746 J

Mass of dry ice required =E/L

= 24*3600*7400*746/573700

= 831380 kg answer


answered by: ANURANJAN SARSAM
Answer #2

Energy required for a day E= 24*3600*7400*746 J

Mass of dry ice required =E/L

= 24*3600*7400*746/573700

= 831380 kg answer